OpenStudy (anonymous):

Help will fan and medal !!:)) Expand the following: A) (y-4)^3x^2

OpenStudy (anonymous):

Oh and btw the x is a variable not a multiplication sign:))

TheSmartOne (thesmartone):

Is it \[(y-4)^{3x^2}\]

OpenStudy (anonymous):

Yes thats what it is @TheSmartOne :))

OpenStudy (acal21):

the problem expanded x^2y^3−12x^2y^2+48x^2y−64x^2

OpenStudy (anonymous):

What are all those diamond question marks?? @acal21 ,Thats confusing..

TheSmartOne (thesmartone):

\[y ^{3x^2}-4^{3x^2}\]

TheSmartOne (thesmartone):

Wouldn't that be the answer @acal21

OpenStudy (anonymous):

Yo @TheSmartOne Thank you so much for taking the time for helping me with your work..It really helped,Have a great night !!:))

TheSmartOne (thesmartone):

You too

OpenStudy (jziggy):

@TheSmartOne do you mind explaining your thought process on this? I'm having trouble seeing how this works.

TheSmartOne (thesmartone):

I thought that you could use distributive property

OpenStudy (jziggy):

(a+b)^n =/= a^n + b^n

OpenStudy (jziggy):

From my attempt to expand it: (y-4)^(3x^2) = ((y-4)^3)^x^2 = [y^3 - 12y^2 +48y - 64]^x^2 which is as far as I think it goes.....